OpenStudy (anonymous):

Tell whether or not the following statement are ture for all the numbers. If the statement is not true, give a counterexample. If a^2=b^2, then a=b. help!!

OpenStudy (anonymous):

aahh.. now this one u hv to tell...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

okk tell me a=2 and b= -2 (-2)^2= 2^2=

OpenStudy (anonymous):

so answer is true or false?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

im guessing its true so the statement is true for all the numbers right?

OpenStudy (anonymous):

uff how its true... is a=b in our example?

OpenStudy (anonymous):

a^2=b^2 but a is not equal to b in our example.... so its false dear... n example i have given you....

OpenStudy (anonymous):

so that would be the counterexample right?

OpenStudy (anonymous):

yeahhh....

OpenStudy (anonymous):

oook thnx